A motor gives a disk A an angular acceleration of aA (2t+ 3) rad/s , where t is in seconds. If the inidial angular velocity of the disk is Dg 5 rads, determine the magnitude of the velocity of biock B when t 2 s aisa v-1.5 m/s ve 15 m/s ve-7 m/s va-2.25 mls
The velocity ofa particle is defined by v (-2i+(3-20A, m/s where t is in seconds. If r 0 when t 0, determine the displacement of the particle during the time interval t As (-8/-4A,m As (-6i+-6A. m As -10/-2A.m As (-4, m 1 s and t 4 s.
Answer #1

duw dt 2 w-5 todr=vdt Saeat t dr t+ t-t2) tIs24+(3c)- lt=4s =-(4)i+(304) -(4))ニー8-4 erasts a-43)-2423) =-6i-6 ti-63m w AS